### v4

# T-araVLN: Translator for Agricultural Robotic Agents on Vision-and-Language Navigation

Abstract
Agricultural robotic agents have been becoming powerful helpers in a wide range of agricultural tasks, however, still heavily rely on manual operation or fixed railways for movement. To address this limitation, the AgriVLN method and the A2A benchmark pioneeringly extend Vision-and-Language Navigation (VLN) to the agricultural domain, enabling agents to navigate to the target positions following the natural language instructions. AgriVLN effectively understands the simple instructions, but often misunderstands the complex ones. To bridge this gap, we propose the method of Translator for Agricultural Robotic Agents on Vision-and-Language Navigation (T-araVLN), in which the Instruction Translator module translates the original instruction to be more refined and precise. When evaluated on the A2A benchmark, our T-araVLN effectively improves Success Rate from 0.47 to 0.63 and reduces Navigation Error from 2.91m to 2.28m, demonstrating the state-of-the-art performance in the agricultural domain. Code: https://github.com/AlexTraveling/T-araVLN.
Index Terms—  Vision-and-Language Navigation, Agricultural Robotic Agent, Large Language Model
1 Introduction
Fig. 1: T-araVLN (right) v.s. Baseline (left) on a simple example: Unspecific instructions such as “close” are translated to representational instructions such as “she occupies all your camera view”, leading to easier alignment between linguistic and visual inputs.
Fig. 2: T-araVLN methodology illustration: Prompted by the five translation principles with the task description and the format restriction, T-araVLN translates the original instruction to be more refined and precise, then understands both linguistic and visual inputs to predict the low-level action sequence. (*Base model’s elaborate prediction process is detailed in AgriVLN [1])
Agricultural robotic agents have been becoming powerful helpers in a wide range of agricultural tasks, such as phenotypic measurement [2], disease detection [3], and maturity recognition [4]. However, most of them still heavily rely on manual operation or fixed railways for movement, resulting in limited mobility and poor adaptability in diverse scenarios.
In contrast, Vision-and-Language Navigation (VLN) enables agents to follow the natural language instructions to navigate to the target positions [5, 6], having demonstrated strong performance across various domains [7], such as R2R [5] for indoor room, TOUCHDOWN [8] for urban street, and AerialVLN [9] for aerial space. Motivated by prior Vision-Language Model-based methods [10, 11, 12], Zhao et al. [1] proposes the AgriVLN method and the A2A benchmark, pioneeringly extending VLN to the agricultural domain. To better align with the practical speaking tone of agricultural workers, A2A’s instructions are deliberately designed with many noises and mistakes, which inevitably brings significant challenges for the alignment between linguistic and visual inputs [13]. The result is that AgriVLN effectively understands the simple instructions, but often misunderstands the complex ones, as illustrated in the left part of Figure 1. This interesting phenomenon brings us a question: Can we polish an instruction to make it easier to understand?
The most relevant study is VLN-Trans [14], which trains the translator module to convert original instructions into easy-to-follow sub-instruction representations, effectively solving two common types of instruction issue. However, the instruction issue types in A2A are much more complicated than in other mainstream VLN benchmarks [5, 15, 16], which makes training-based methods like VLN-Trans [14] and MTST [17] struggling with the circumstances out of training data, leaving a large gap on generalization capability.
To bridge this gap, we propose the method of Translator for Agricultural Robotic Agents on Vision-and-Language Navigation (T-araVLN), as illustrated in Figure 1. First, we propose the module of Instruction Translator, which translates the original instruction to be more refined and precise. Second, we integrate the Instruction Translator module into the base model of AgriVLN, to establish our T-araVLN method. Third, we evaluate T-araVLN on the A2A benchmark, effectively improving Success Rate from 0.47 to 0.63 and reducing Navigation Error from 2.91m to 2.28m, which achieves the state-of-the-art performance in the agricultural VLN domain. We share a simple example in Figure 1 to illustrate the difference between T-araVLN and Baseline.
In summary, our main contributions are as follows:
  • Instruction Translator, a Large Language Model-based module, translates instructions from noisy and mistaken to refined and precise.
  • T-araVLN, a VLN method integrating Instruction Translator, navigates agricultural robots to the target positions following the natural language instructions.
  • We implement Ablation Experiment proving the effectiveness of Instruction Translator and Comparison Experiment proving the SOTA performance of T-araVLN.
Fig. 3: Qualitative Experiment results on the representative episode: The original instruction includes several noises and mistakes, leading to the challenging alignment between linguistic and visual inputs. Our T-araVLN, however, translates the instruction to be refined and precise, effectively improving the alignment performance. ( represents correct prediction after rational reasoning; represents correct prediction after illogical reasoning; represents wrong prediction after illogical reasoning. Only for a complete demonstration, the baseline model is reset to the ground-truth path after deviation. Zoom in for better observation.)
2 Methodology
2.1 Task Definition
The task of Agricultural Vision-and-Language Navigation [1] is defined as follows: In each episode, the model is given an instruction in natural language, denoted as $W=\langle w_{1},w_{2},\dots,w_{L}\rangle$, where $L$ is the number of words. At each time step $t$, the model is given the front-facing RGB image $I_{t}$. The purpose is understanding both $W$ and $I_{t}$, to select the best low-level action $\hat{a_{t}}$ from action space $\{\texttt{FORWARD}$, LEFT ROTATE, RIGHT ROTATE, $\texttt{STOP}\}$, thereby leading the robot navigate from the starting point to the destination.
2.2 T-araVLN
We propose the method of Translator for Agricultural Robotic Agents on Vision-and-Language Navigation (T-araVLN), as illustrated in Figure 2.
2.2.1 Instruction Translator
We utilize Large Language Model (LLM) to establish the module of Instruction Translator, denoted as $\mathcal{T}$. For a raw instruction $W$ in casual spoken-language style containing complete semantics, $\mathcal{T}$ translates it to formal written-language style retaining only essential semantics, denoted as $W^{\prime}$, defined as:
$R,W^{\prime}=\mathcal{T}\>(W\;\|\;P_{\mathcal{T}})$ (1)
where $R$ denotes the explicit reasoning process. $\|$ denotes the concatenation. $P_{\mathcal{T}}$ denotes the prompt template for $\mathcal{T}$, defined as:
$P_{\mathcal{T}}=p_{\text{description}}\;\|\;\mathop{\|}_{i=1}^{K}(p_{i},e_{i})\;\|\;p_{\text{format}}$ (2)
where $K\in\{1,\dots,5\}$. $p_{\text{description}}$ denotes the task description. $p_{\text{format}}$ denotes the format restriction on input and output. $\{p_{1},\dots,p_{5}\}$ denote the five translation principles: Inessential Removal, Errors Revision, High-Low Separation, Representational Rotation, and Representational Movement, discussed in detail in Figure 2. Meanwhile, we design every $p_{i}$ in one-shot format with one example denoted as $e_{i}$. After being translated, $W^{\prime}$ is easier than $W$ to understand for the downstream VLN decision-making model.
2.2.2 Base Model
We follow the architecture of AgriVLN [1] as our base model, denoted as $\mathcal{D}$, in which the module of $\mathcal{T}$ is integrated between the layer of instruction input and the module of Subtask List. At each time step $t$, $\mathcal{D}$ attempts understanding both translated instruction $W^{\prime}$ and camera image $I_{t}$, to select the best low-level action $\hat{a_{t}}$ for navigation, defined as:
$\langle\hat{a_{1}},\hat{a_{2}},\dots,\hat{a_{t^{\prime}}}\rangle=\mathcal{D}\>(W^{\prime},\langle I_{1},I_{2},\dots,I_{t^{\prime}}\rangle)$ (3)
where $t^{\prime}$ denotes the ending time step, at which at least one of the following conditions happened: 1. $\hat{a_{t^{\prime}}}$ = STOP; 2. predicted sequence $\langle\hat{a_{t^{\prime}-\tau}},\hat{a_{t^{\prime}-\tau+1}},\dots,\hat{a_{t^{\prime}}}\rangle$ is deviated to label sequence $\langle a_{t^{\prime}-\tau},a_{t^{\prime}-\tau+1},\dots,a_{t^{\prime}}\rangle$, where $\tau$ denotes the time step threshold [^1] More details about deviation judgement are available in AgriVLN [1].; 3. $t^{\prime}$ reaches the upper limitation.
2.3 Evaluation Metrics
We follow two standard evaluation metrics for VLN: Success Rate (SR) and Navigation Error (NE) [5]. In addition, we introduce BERTScoreF1 [18] to calculate the semantic similarity between the original instruction and the translated one, in which a lower score indicates a greater degree of modification, leading to a more aggressive translation preference; conversely, a higher score indicates a smaller degree of modification, leading to a more conservative translation preference.
3 Experiments
3.1 Experimental Settings
We implement all experiments on the A2A [1] benchmark, and access all Large Language Models (LLMs) through API.
3.2 Qualitative Experiment
Toward better understanding for readers, we implement Qualitative Experiment on the representative episode, as illustrated and discussed in Figure 3.
# Method Translator A2A ( low-complexity ) A2A ( high-complexity ) A2A
[…]
1 Random - 0.13 7.30 - 0.04 6.74 - 0.09 7.03 -
2 Fixed - 0.00 0.00 - 0.06 6.32 - 0.03 3.06 -
State-of-the-Art
3 SIA-VLN [19] - 0.52 1.46 - 0.08 5.12 - 0.31 3.24 -
4 DILLM-VLN [20] - 0.41 1.36 - 0.32 3.90 - 0.36 2.60 -
5 AgriVLN [1] - 0.58 2.32 - 0.35 3.54 - 0.47 2.91 -
Ours
6 T-araVLNc Claude-3.7-sonnet prompted 0.66 2.12 0.9096 0.49 3.88 0.9112 0.58 2.97 0.9105
7 T-araVLNd DeepSeek-r1 prompted 0.70 1.68 0.8977 0.47 3.49 0.8982 0.59 2.55 0.8980
8 T-araVLNg GPT-4.1 prompted 0.80 1.41 0.9028 0.46 3.22 0.9039 0.63 2.28 0.9035
9 Human - 0.93 0.32 - 0.80 0.82 - 0.87 0.57 -
Table 1: Comparison Experiment results between T-araVLN and state-of-the-art methods on the low-complexity portion (subtask == 2), the high-complexity portion (subtask $\geq$ 3), and the whole of A2A. Bold and underline represent the best and worst scores, respectively.
3.3 Comparison Experiment
We compare T-araVLN with three existing state-of-the-art models: SIA-VLN [19], DILLM-VLN [20] and AgriVLN [1]. In addition, the methods of Random, Fixed and Human are reproduced as the lower and upper bounds, respectively.
The results of Comparison Experiment are shown in Table 1. On the whole of A2A, our T-araVLN with GPT-4.1-based translator (#8) achieves SR of 0.63 and NE of 2.28, surpassing all the three baselines (#3, #4 and #5) on both SR and NE. On the low-complexity portion, our T-araVLN even substantially improves SR from 0.58 to 0.80, which nearly approaches the performance of human. In summary, Comparison Experiment proves the significant effectiveness of the Instruction Translator module, which leads to the state-of-the-art performance of T-araVLN in the agricultural VLN domain.
3.4 Ablation Experiment
3.4.1 Large Language Model
We test T-araVLN with two additional LLMs of comparable size and performance, as shown in Table 1. When the Instruction Translator module is driven by Claude-3.7-sonnet (#6) or DeepSeek-r1 (#7) [21], the BERTScores are 0.9105 or 0.8980, demonstrating more conservative or more radical translation strategies, respectively. However, SR decreases by 5 or 4 percentage points and NE increases by 0.69 or 0.27 meters, respectively, which suggests that overly conservation may bring tedious noises while overly radical may delete essential semantics. Compared to them, GPT-4.1 achieves the best trade-off. Therefore, we select GPT-4.1 as the LLM for the Instruction Translator module in T-araVLN.
3.4.2 Translation Principle
We ablate the five translation principles in the module of Instruction Translator, in which we follow GPT-4.1 as LLM, as shown in Table 2 and Figure 4. When every single principle is applied (#11, #12, #13, #14 or #15), the performance improves by 2 $\sim$ 10 percentage points on SR and 0.01 $\sim$ 0.52 meters on NE, which indicates the effectiveness of every principle. When all principles are combined (#16), the performance reaches the best, i.e., our proposed T-araVLN, which indicates that every principle contributes to a different emphasis of translation, meanwhile, all principles collaborate without significant interference.
[…]
10 Baseline 0.47 2.91 -
11    + Inessential Removal 0.50 2.80 0.9041
12    + Errors Revision 0.51 2.85 0.9028
13    + High-Low Separation 0.50 2.90 0.9052
14    + Representational Rotation 0.49 2.80 0.9009
15    + Representational Movement 0.57 2.39 0.9005
16 T-araVLN (Ours) 0.63 2.28 0.9035
Table 2: Ablation Experiment results on the five translation principles in Instruction Translator. “+” represents integration. Bold and underline represent the best and worst scores, respectively.
(a) Success Rate $\uparrow$
(b) Navigation Error $\downarrow$
Fig. 4: Ablation Experiment illustration on the five translation principles in Instruction Translator. “+” represents integration.
4 Conclusion
In this paper, we propose the Translator for Agricultural Robotic Agents on Vision-and-Language Navigation (T-araVLN) method, in which the Instruction Translator module translates the original instruction to be more refined and precise. When evaluated on the A2A benchmark, our T-araVLN effectively improves Success Rate from 0.47 to 0.63 and reduces Navigation Error from 2.91m to 2.28m, demonstrating the state-of-the-art performance in the agricultural domain.
Acknowledgment
This study is supported by the National Natural Science Foundation of China’s Research on Distributed Real-Time Complex Event Processing for Intelligent Greenhouse Internet of Things (Grant No. 61601471).